OpenStudy (anonymous):

how do i do this problem Danny cuts a 19 1/2meters long ribbon into smaller pieces. Each piece is 3/4meters long. Danny wants to distribute the ribbon pieces equally among his friends. Which of the following expressions should he use to find out the total number of friends he can distribute the ribbon pieces among?

OpenStudy (anonymous):

the total number of friends he can distribute the ribbon pieces among = length of the ribbon/ length of each piece \[= \frac{ 19 \frac{1}{2} }{\frac{3}{4} }\]

OpenStudy (anonymous):

oh ok so i divide it... ok thanks i thought so but i wasnt sure

OpenStudy (eyust707):

Remember that \(19 {1 \over 2}= {39 \over 2}\)

OpenStudy (anonymous):

@popie You are welcome:)